# Asymptotic Normality of Subgraph Counts in Sparse Inhomogeneous Random Graphs

### v1

@@ 1. Introduction
A graphon is a symmetric measurable function $W:[0,1]^{2}\to[0,1]$, that is, $W(x,y)=W(y,x)$ for all $x,y\in[0,1]$. Graphons arise as limit objects of sequences of large graphs and have attracted significant attention in recent years. They form a powerful bridge between combinatorics and analysis and have found applications across a range of disciplines, including statistical physics, probability, and statistics; see, for example, [13, 15, 4, 19, 20]. For a comprehensive exposition of the theory of graph limits, we refer to Lovász [41]. Graphons also provide a natural probabilistic model for generating inhomogeneous random graphs, which generalize the classical (homogeneous) Erdős–Rényi random graph, where edge probabilities vary across vertex pairs according to the underlying graphon (see [12, 24, 42, 9] among others).
Random graphs generated from a graphon model produces dense graphs, which have $O(n^{2})$ edges with high probability. In contrast, many real-world networks, arising from social, biological, and technological applications, are observed to be sparse. To model such sparsity, the sparse graphon model has been proposed, which introduces a sparsity parameter $\rho_{n}$ for controlling the edge density (see [8, 7, 36, 43] references therein):
@@
Given a sparsity parameter $\rho_{n}\in(0,1)$ and a graphon $W$ with $\int_{[0,1]^{2}}W(x,y)\mathrm{d}x\mathrm{d}y>0$, the $W$-random graph with sparsity $\rho_{n}$ on the vertex set $n]:=\{1,2,\ldots,n\}$, hereafter denoted by $G(n,\rho_{n},W)$, is obtained by connecting the vertices $i$ and $j$ with probability $\rho_{n}W(U_{i},U_{j})$, independently for all $1\leq i<j\leq n$, where $\{U_{i}:1\leq i\leq n\}$ is a sequence of i.i.d. $U[0,1]$ random variables.
The model in Definition 1.1 encompasses many well-known network models, including the classical Erdős-Rényi random graph model (where $W$ is a constant function), the stochastic block model [7, 31] (where $W$ is a block function), latent space models [29], random dot-product graphs [1, 49], and random geometric graphs [47], among others. Over the years, the sparse graphon model has emerged as a fundamental tool in modern network analysis, with applications in community detection, subgraph count statistics, and nonparametric estimation of graph parameters, among others.
Among the basic features of a network are its subgraph (motif) counts, that is, the frequencies of particular patterns (subgraphs), such as the number or density of edges, triangles, or stars within the network. These encode important structural information about the geometry of a network, and many features of practical interest can be derived from motif counts, such as the clustering coefficient [55], degree distribution [48], and transitivity [30] (see [50] for others). In the framework of the graphon model, subgraph counts are often referred to as network moments [8, 14], which provide an important tool for inferring properties of the underlying graphon (see [6, 38, 54, 27, 43, 45] and references therein). Consequently, understanding the asymptotic properties of subgraph counts in $W$-random graphs is a problem of central importance in statistical network analysis. To this end, given a fixed graph $H=(V(H),E(H))$ denote by $X_{n}(H,G_{n})$ the number of copies of $H$ in the $W$-random graph $G(n,\rho_{n},W)$. More formally,
@@
where, for any set $S\subseteq[n]$, $\mathscr{G}_{H}(S)$ denotes the collection of all subgraphs of the complete graph $K_{|S|}$ on the vertex set $S$ which are isomorphic to HH. [^1] Note that we count unlabelled copies of $H$. Several other papers count labelled copies, where $X(H,G_{n})$ is multiplied by $|\text{Aut}(H)|$. Note that
@@
$\mathbb{E}[X(H,G_{n})]$ $=\sum_{1\leq i_{1}<\cdots<i_{|V(H)|}\leq n}\sum_{H^{\prime}\in\mathscr{G}_{H}(\{i_{1},\ldots,i_{|V(H)|}\})}\prod_{(i_{s},i_{t})\in E(H^{\prime})}\rho_{n}\mathbb{E}[W(U_{i_{s}}U_{i_{t}})]$
@@
Classical results of Ruciński [51], Barbour et al. [3], Nowicki [46], Janson [33] provide a complete characterization of the asymptotic normality of $X(H,G_{n})$ (after appropriate centering and scaling) when $G_{n}\sim G(n,\rho_{n})$ is the (homogeneous) Erdős-Rényi model (see [34, Chapter 6] for comprehensive treatment). For $G_{n}\sim G(n,\rho_{n},W)$ sampled from the sparse random graphon model, the fluctuation of $X(H,G_{n})$ was first studied in the seminal paper of Bickel et al. [8]. This result has been refined and extended in many directions (see the discussion Section 2.4). However, existing results remain incomplete: they apply only to specific types of subgraphs and often do not cover the full sparsity regime. Our objective this paper is to fill these missing gaps. Specifically, we establish the asymptotic normality of $Z(H,G_{n})$ for all fixed graphs $H$ in the entire allowable sparsity regime (that is, whenever $\rho_{n}$ is above the containment threshold of $H$ in $G(n,\rho_{n},W)$). In fact, we provide a more fine-grained analysis that decouples the randomness of the vertex labels from the randomness of the edges and identifies distinct regimes of sparsity where the two components contribute, either separately or simultaneously, to the fluctuations of $Z(H,G_{n})$. This relies on decomposing the centered subgraph count as follows:
$\Delta(H,G_{n})$ $:=X(H,G_{n})-\mathbb{E}[X(H,G_{n})]$
$=\big(X(H,G_{n})-\mathbb{E}[X(H,G_{n})\mid\mathcal{F}(\bm{U}_{n})]\big)+\big(\mathbb{E}[X(H,G_{n})\mid\mathcal{F}(\bm{U}_{n})]-\mathbb{E}[X(H,G_{n})]\big)$
@@
where $\mathcal{F}(\bm{U}_{n})$ is the sigma-algebra generated by the collection $\bm{U}_{n}=\{U_{1},U_{2},\ldots,U_{n}\}$. Note that, conditioned on $\bm{U}_{n}$, the randomness in the first term $\Delta_{1}(H,G_{n})$ arises solely from the edges of $G_{n}$. On the other hand, $\Delta_{2}(H,G_{n})$ depends only on vertex labels $\{U_{1},U_{2},\ldots,U_{n}\}$, that is, it is $\mathcal{F}(\bm{U}_{n})$-measurable. Contributions of $\Delta_{1}(H,G_{n})$ and $\Delta_{2}(H,G_{n})$ to the total variance of $X(H,G_{n})$ depend on the regime of sparsity and also on a notion of $H$-regularity of the graphon $W$ (see Definition 2.4). Specifically, we obtain the following results:
  • If $W$ is $H$-regular, the contribution of $\Delta_{2}(H,G_{n})$ to the variance of $X(H,G_{n})$ is asymptotically negligible, and the limiting distribution of $X(H,G_{n})$ is governed solely by $\Delta_{1}(H,G_{n})$ (see Theorem 2.6).
  • If $W$ is not $H$-regular, the asymptotic behaviors of $\Delta_{1}(H,G_{n})$ and $\Delta_{2}(H,G_{n})$ exhibits a phase transition. Below a critical threshold, $\Delta_{2}(H,G_{n})$ contributes negligibly to the variance of $X(H,G_{n})$, and the fluctuations are driven entirely by $\Delta_{1}(H,G_{n})$. Above the threshold, the roles reverse, with $\Delta_{1}(H,G_{n})$ becoming negligible and $\Delta_{2}(H,G_{n})$ governing the fluctuations. At the threshold, both terms contribute, and the appropriately standardized pair $(\Delta_{1}(H,G_{n}),\Delta_{2}(H,G_{n}))$ converges jointly to independent normal distributions with nontrivial variances (see Theorem 2.11).
The above result combined show that for any subgraph $H=(V(H),E(H))$, with $|E(H)|\geq 1$, $Z(H,G_{n})$ is asymptotically normal whenever $\rho_{n}$ is above the containment threshold of $H$ in $G(n,W)$ (see Corollary 2.13). This completes the program, initiated Bickel et al. [8], of deriving the asymptotic distribution of subgraph counts (network moments) in sparse graphon models in full generality. In particular, it also answers a question posed by Hladkỳ et al. [28] about the distribution of the number of cliques. As a byproduct of our analysis we obtain a conditional central limit theorem for $\Delta_{1}(H,G_{n})$ (in the Wasserstein distance) that might be of independent interest (see Proposition 2.15).
@@
Throughout the paper we will use the following asymptotic notations: For two sequences $a_{n}$ and $b_{n}$ we will write $a_{n}\lesssim b_{n}$ and $a_{n}=O(b_{n})$ if for all $n$ large enough $a_{n}\leq C_{1}b_{n}$, for some constant $C_{1}>0$. Similarly, $a_{n}\gtrsim b_{n}$ will mean $a_{n}\geq C_{2}b_{n}$ and $a_{n}\asymp b_{n}$ will mean $C_{2}b_{n}\leq a_{n}\leq C_{1}b_{n}$, for $n$ large enough and constants $C_{1},C_{2}>0$. Subscripts in the above notation, for example $\lesssim_{\square}$, $O_{\square}$, $\gtrsim_{\square}$, and $\asymp_{\square}$ denote that the hidden constants may depend on the subscripted parameters. Moreover, $a_{n}\ll b_{n}$, $a_{n}\gg b_{n}$, and $a_{n}\sim b_{n}$ will mean $a_{n}/b_{n}\rightarrow 0$, $a_{n}/b_{n}\rightarrow\infty$, $a_{n}/b_{n}\rightarrow 1$, as $n\rightarrow\infty$.
@@ 2. Fluctuations of Subgraph Counts in Sparse $W$-Random Graphs
In this section we will state our main result on the asymptotic distribution $Z(H,G_{n})$, where $G_{n}\sim G(n,\rho_{n},W)$ is a $W$-random graph with sparsity $\rho_{n}$. The section is organized as follows.
@@ 2.1. Preliminaries
The preliminary towards deriving the asymptotic distribution is $Z(H,G_{n})$ is to figure out rate at which $\rho_{n}$ has to diverge such that $G_{n}\sim G(n,\rho_{n},W)$ contains at least one (or infinitely many) copies of $H$. For the classical Erdős-Rényi model this is standard fare in the random graphs literature (see [34, Chapter 3]). In particular, it depends on whether or not the graph $H$ is balanced:
@@
[34, Equation (3.6)] For a fixed graph $H=(V(H),E(H))$, define
@@
where the maximum is taken over all possible non-empty subgraphs of $H.$ If the maximum is attained by $H$ itself, that is, $m(H)=\frac{|E(H)|}{|V(H)|}$, then $H$ said to be a balanced graph. Otherwise, $H$ said to be an unbalanced graph. Moreover, if the maximum in (2.1) uniquely attained $H$, then $H$ is called a strictly balanced graph.
@@
It is well-known from classical results [25, 10] that for an Erdős-Rényi random $G(n,\rho_{n})$, the threshold for the appearance of a fixed graph $H=(V(H),E(H))$ is determined by the condition $n\rho_{n}^{m(H)}\asymp 1$ (see [34, Theorem 3.4]). The proof of this result can be easily adapted to show that the threshold remains the same for the inhomogeneous random graph model $G(n,\rho,W)$ as well. We summarize this result in the following proposition. A proof is included in Section 3 for the sake of completeness.
@@
$X(H,G_{n})\stackrel{{\scriptstyle P}}{{\rightarrow}}\left\{\begin{array}[]{ll}0&\text{if }n\rho_{n}^{m(H)}\ll 1,\\ \infty&\text{if }n\rho_{n}^{m(H)}\gg 1.\end{array}\right$.
@@ 2.2. Statements of the Results
$Z(H,G_{n})=\frac{X(H,G_{n})-\mathbb{E}[X(H,G_{n})]}{\sqrt{\mathrm{Var}[X(H,G_{n})]}}=\frac{\Delta(H,G_{n})}{\mathrm{Var}[\Delta(H,G_{n})]},$ (2.2)
where $\Delta(H,G_{n}):=X(H,G_{n})-\mathbb{E}[X(H,G_{n})]$. Recalling the the decomposition in (2.11), note that
@@
A graphon $W$ is said to be $H$-regular if, for almost every $x\in[0,\,1]$,
@@
for all $a\in V(H)$. Hence, if $\overline{t}(x,H,W)$ is a constant almost everywhere, then the constant must be $t(H,W)$. Hence, in other words, a graphon $W$ is $H$-regular if the homomorphism density of $H$ in $W$ when one of the vertices of $H$ is marked, is a constant independent of the value of the marking. Alternatively, $H$-regularity is equivalent to the condition that the $U$-statistic $\Delta_{2}(H,\,G_{n})$ is first-order degenerate.
@@
Hence, the notion of $K_{2}$-regularity coincides with the standard notion of degree regularity, where the degree function $d_{W}(x):=\int_{0}^{1}W(x,y)\mathrm{d}y$ is constant almost everywhere.
In the dense regime (where $\rho_{n}\asymp 1$), [5, 28] showed that $H$-regularity can lead $X(H,G_{n})$ to have non-Gaussian fluctuations. The situation, however, is very different in the sparse regime. In this case, $H$-regularity implies that $\Delta_{2}(H,G_{n})$ has asymptotically negligible contribution to the total variance of $X(H,G_{n})$. Consequently, the asymptotic distribution of $X(H,G_{n})$ is determined entirely by the fluctuations of $\Delta_{1}(H,G_{n})$.
@@
Suppose Assumption 2.3 holds and $W:[0,1]^{2}\rightarrow\mathbb{R}_{\geq 0}$ is a $H$-regular graphon. Then,
@@
$Z(H,G_{n})=\frac{\Delta(H,G_{n})}{\sqrt{\mathrm{Var}[\Delta(H,G_{n})]}}\stackrel{{\scriptstyle D}}{{\rightarrow}}N(0,\,1),$ (2.6)
@@
Next, we consider the case where $W$ is not $H$-regular. In this case, the situation is more delicate. Specifically, an additional threshold on $\rho_{n}$ emerges, which determines whether or not $\Delta_{2}(H,G_{n})$ contributes to the fluctuations of $X(H,G_{n})$. To describe this threshold we need the following definition:
@@
[34, Equation (3.17)] For a fixed graph $H=(V(H),E(H))$ define
@@
Cycles, trees and complete graphs are strongly balanced. Figure 2 (b) is an example of graph $H$ which is not strongly balanced. In this case, for the subgraph $F$ marked by the red dotted circle
$\frac{|E(F)|}{|V(F)|-1}=\frac{3}{2}>\frac{4}{3}=\frac{|E(H)|}{|V(H)|-1}.$
@@
The notion of a strongly balanced graph was introduced by [53], where it is shown that $n\rho_{n}^{m_{1}(H)}\asymp 1$ is the threshold for the containment of a rooted copy of $H$ in the Erdős-Rényi model $G(n,\rho_{n})$. Recently, Maugis [44] showed that for the sparse graphon model the number of rooted copies of $H$ (appropriately standardized) has Gaussian fluctuations, whenever $n\rho_{n}^{m_{1}(H)}\gg 1$. In Theorem 2.11 we prove that when $W$ is not $H$-regular, then $n\rho_{n}^{m_{1}(H)}\asymp 1$ marks the threshold where the asymptotic behavior of $\Delta_{1}(H,G_{n})$ and $\Delta_{2}(H,G_{n})$ undergoes a phase transition:
  • when $n\rho_{n}^{m_{1}(H)}\ll 1$, $\Delta_{1}(H,G_{n})$ is asymptotically Gaussian and $\Delta_{2}(H,G_{n})$ is asymptotically negligible;
  • when $n\rho_{n}^{m_{1}(H)}\gg 1$, $\Delta_{2}(H,G_{n})$ is asymptotically Gaussian and $\Delta_{1}(H,G_{n})$ is asymptotically negligible; and
  • when $n\rho_{n}^{m_{1}(H)}\asymp 1$, $(\Delta_{1}(H,G_{n}),\Delta_{2}(H,G_{n}))$ are asymptotically jointly Gaussian.

### v2

# arXiv:2512.12937v2[math.PR] 01 Mar 2026

@@ 1. Introduction
A graphon is a symmetric measurable function $W:[0,1]^{2}\to[0,1]$, that is, $W(x,y)=W(y,x)$ for all $x,y\in[0,1]$. Graphons arise as limit objects of sequences of large graphs and have attracted significant attention in recent years. They form a powerful bridge between combinatorics and analysis and have found applications across a range of disciplines, including statistical physics, probability, and statistics; see, for example, Borgs et al. (2008, 2012), Basak and Mukherjee (2017), Chatterjee and Diaconis (2013), Chatterjee and Varadhan (2011). For a comprehensive exposition of the theory of graph limits, we refer to Lovász (2012). Graphons also provide a natural probabilistic model for generating inhomogeneous random graphs, which generalize the classical (homogeneous) Erdős–Rényi random graph, where edge probabilities vary across vertex pairs according to the underlying graphon (see Bollobás et al. (2007), Diaconis and Freedman (1981), Lovász and Szegedy (2006), Boguná and Pastor-Satorras (2003) among others).
Random graphs generated from a graphon model produces dense graphs, which have $O(n^{2})$ edges with high probability. In contrast, many real-world networks, arising from social, biological, and technological applications, are observed to be sparse. To model such networks, the sparse graphon model has been proposed, which introduces a parameter $\rho_{n}$ to control the edge density (see Bickel et al. (2011), Bickel and Chen (2009), Klopp et al. (2017), Lunde and Sarkar (2023) references therein):
@@
Given a sparsity parameter $\rho_{n}\in(0,1)$ and a graphon $W$ with $\int_{[0,1]^{2}}W(x,y)\mathrm{d}x\mathrm{d}y>0$, the $W$-random graph with sparsity $\rho_{n}$ on the vertex set $n]:=\{1,2,\ldots,n\}$, hereafter denoted by $G(n,\rho_{n},W)$, is obtained by connecting the vertices $i$ and $j$ with probability $\rho_{n}W(U_{i},U_{j})$, independently for all $1\leq i<j\leq n$, where $\{U_{i}:1\leq i\leq n\}$ is a sequence of i.i.d. $U([0,1])$ random variables.
The model in Definition 1.1 encompasses many well-known network models, including the classical Erdős-Rényi random graph model (where $W$ is a constant function), the stochastic block model (Bickel and Chen, 2009, Holland et al., 1983) (where $W$ is a block function), latent space models (Hoff et al., 2002), random dot-product graphs (Athreya et al., 2018, Rubin-Delanchy et al., 2022), and random geometric graphs (Penrose, 2003), among others. Over the years, the sparse graphon model has emerged as a fundamental tool in modern network analysis, with applications in community detection, subgraph count statistics, and nonparametric estimation of graph parameters, among others.
Among the basic features of a network are its subgraph (motif) counts, that is, the frequencies of particular patterns (subgraphs), such as the number or density of edges, triangles, or stars within the network. These encode important structural information about the geometry of a network, and many features of practical interest can be derived from motif counts, such as the clustering coefficient (Watts and Strogatz, 1998), degree distribution (Pržulj, 2007), and transitivity (Holland and Leinhardt, 1971) (see Rubinov and Sporns (2010) for others). In the framework of the graphon model, subgraph counts are often referred to as network moments (Bickel et al., 2011, Borgs et al., 2010), which provide an important tool for inferring properties of the underlying graphon (see Bhattacharyya and Bickel (2015), Lin et al. (2022), Shao et al. (2022), Green and Shalizi (2022), Lunde and Sarkar (2023), Maugis et al. (2020) and references therein). Consequently, understanding the asymptotic properties of subgraph counts in $W$-random graphs is a problem of central importance in statistical network analysis. To this end, given a fixed graph $H=(V(H),E(H))$ denote by $X_{n}(H,G_{n})$ the number of copies of $H$ in the $W$-random graph $G_{n}\sim G(n,\rho_{n},W)$. More formally,
@@
where, $a_{ij}=1$, if $(i,j)$ is an edge in $G_{n}$ and 0 otherwise, for $1\leq i<j\leq n$, and, for any set $S\subseteq[n]$, $\mathscr{G}_{H}(S)$ denotes the collection of all subgraphs of the complete graph $K_{|S|}$ on the vertex set $S$ which are isomorphic to HH. [^1] Note that we count unlabelled copies of $H$. Several other papers count labelled copies, where $X_{n}(H,G_{n})$ is multiplied by $|\text{Aut}(H)|$. Note that
@@
$\mathbb{E}[X_{n}(H,G_{n})]$ $=\sum_{1\leq i_{1}<\cdots<i_{|V(H)|}\leq n}\sum_{H^{\prime}\in\mathscr{G}_{H}(\{i_{1},\ldots,i_{|V(H)|}\})}\mathbb{E}\left[\prod_{(i_{s},i_{t})\in E(H^{\prime})}[\rho_{n}W(U_{i_{s}}U_{i_{t}})]\right]$
@@
Classical results of Ruciński (1988), Barbour et al. (1989), Nowicki (1989), Janson (1990) provide a complete characterization of the asymptotic normality of $X_{n}(H,G_{n})$ (after appropriate centering and scaling) when $G_{n}\sim G(n,\rho_{n})$ is the (homogeneous) Erdős-Rényi model (see Janson et al. (2011, Chapter 6) for comprehensive treatment). For $G_{n}\sim G(n,\rho_{n},W)$ sampled from the sparse random graphon model, the fluctuation of $X_{n}(H,G_{n})$ was first studied in the seminal paper of Bickel et al. (2011). This result has been refined and extended in many directions (see the discussion Section 2.4). However, existing results remain incomplete: they apply only to specific types of subgraphs and often do not cover the full sparsity regime. Our objective in this paper is to fill these missing gaps. Specifically, we establish the asymptotic normality of $Z_{n}(H,G_{n}),$ the standardized version of $X_{n}(H,G_{n}),$ for all fixed graphs $H$ in the entire allowable sparsity regime (that is, whenever $\rho_{n}$ is above the containment threshold of $H$ in $G(n,\rho_{n},W)$). In fact, we provide a more fine-grained analysis that decouples the randomness of the vertex labels from the randomness of the edges and identifies distinct regimes of sparsity where the two components contribute, either separately or simultaneously, to the fluctuations of $Z_{n}(H,G_{n})$. This relies on decomposing the centered subgraph count as follows:
$\Delta(H,G_{n})$ $:=X_{n}(H,G_{n})-\mathbb{E}[X_{n}(H,G_{n})]$
$=\big(X_{n}(H,G_{n})-\mathbb{E}[X_{n}(H,G_{n})\mid\mathcal{F}(\bm{U}_{n})]\big)+\big(\mathbb{E}[X_{n}(H,G_{n})\mid\mathcal{F}(\bm{U}_{n})]-\mathbb{E}[X_{n}(H,G_{n})]\big)$
@@
where $\mathcal{F}(\bm{U}_{n})$ is the sigma-algebra generated by the collection $\bm{U}_{n}=\{U_{1},U_{2},\ldots,U_{n}\}$. Note that, conditioned on $\bm{U}_{n}$, the randomness in the first term $\Delta_{1}(H,G_{n})$ arises solely from the edges of $G_{n}$. On the other hand, $\Delta_{2}(H,G_{n})$ depends only on vertex labels $\{U_{1},U_{2},\ldots,U_{n}\}$, that is, it is $\mathcal{F}(\bm{U}_{n})$-measurable. Contributions of $\Delta_{1}(H,G_{n})$ and $\Delta_{2}(H,G_{n})$ to the total variance of $X_{n}(H,G_{n})$ depend on the regime of sparsity and also on a notion of $H$-regularity of the graphon $W$ (see Definition 2.4). Specifically, we obtain the following results:
  • If $W$ is $H$-regular, the contribution of $\Delta_{2}(H,G_{n})$ to the variance of $X_{n}(H,G_{n})$ is asymptotically negligible, and the limiting distribution of $X_{n}(H,G_{n})$ is governed solely by $\Delta_{1}(H,G_{n})$ (see Theorem 2.6).
  • If $W$ is not $H$-regular, the asymptotic behaviors of $\Delta_{1}(H,G_{n})$ and $\Delta_{2}(H,G_{n})$ exhibits a phase transition. Below a critical threshold, $\Delta_{2}(H,G_{n})$ contributes negligibly to the variance of $X_{n}(H,G_{n})$, and the fluctuations are driven entirely by $\Delta_{1}(H,G_{n})$. Above the threshold, the roles reverse, with $\Delta_{1}(H,G_{n})$ becoming negligible and $\Delta_{2}(H,G_{n})$ governing the fluctuations. At the threshold, both terms contribute, and the appropriately standardized pair $(\Delta_{1}(H,G_{n}),\Delta_{2}(H,G_{n}))$ converges jointly to independent normal distributions with nontrivial variances (see Theorem 2.11).
The above results combined show that for any subgraph $H=(V(H),E(H))$, with $|E(H)|\geq 1$, $Z_{n}(H,G_{n})$ is asymptotically normal whenever $\rho_{n}$ is above the containment threshold of $H$ in $G(n,W)$ (see Corollary 2.13). This completes the program, initiated in Bickel et al. (2011), of deriving the asymptotic distribution of subgraph counts (network moments) in sparse graphon models in full generality. In particular, it also answers a question posed by Hladkỳ et al. (2021) about the distribution of the number of cliques. As a byproduct of our analysis we obtain a conditional central limit theorem for $\Delta_{1}(H,G_{n})$ (in the Wasserstein distance) that might be of independent interest (see Proposition 2.15).
@@
Throughout the paper we will use the following asymptotic notations: For two sequences $a_{n}$ and $b_{n}$ we will write $a_{n}\lesssim b_{n}$ and $a_{n}=O(b_{n})$ if for all $n$ large enough $a_{n}\leq C_{1}b_{n}$, for some constant $C_{1}>0$. Similarly, $a_{n}\gtrsim b_{n}$ will mean $a_{n}\geq C_{2}b_{n}.$ Also, $a_{n}\asymp b_{n}$ and $a_{n}=\Theta(b_{n})$ will mean $C_{2}b_{n}\leq a_{n}\leq C_{1}b_{n}$, for $n$ large enough and constants $C_{1},C_{2}>0$. Subscripts in the above notation, for example $\lesssim_{\square}$, $\gtrsim_{\square}$, $O_{\square}$, and $\asymp_{\square}$ denote that the hidden constants may depend on the subscripted parameters. Moreover, $a_{n}\ll b_{n}$, $a_{n}\gg b_{n}$, and $a_{n}\sim b_{n}$ will mean $a_{n}/b_{n}\rightarrow 0$, $a_{n}/b_{n}\rightarrow\infty$, $a_{n}/b_{n}\rightarrow 1$, as $n\rightarrow\infty$.
@@ 2. Fluctuations of Subgraph Counts in Sparse $W$-Random Graphs
In this section we will state our main result on the asymptotic distribution $Z_{n}(H,G_{n})$, where $G_{n}\sim G(n,\rho_{n},W)$ is a $W$-random graph with sparsity $\rho_{n}$.
@@ 2.1. Preliminaries
The preliminary step towards deriving the asymptotic distribution of $Z_{n}(H,G_{n})$ is to figure out the rate at which $\rho_{n}$ has to diverge such that $G_{n}\sim G(n,\rho_{n},W)$ contains at least one (or infinitely many) copies of $H$. For the classical Erdős-Rényi model this is standard fare in the random graphs literature (see Janson et al. (2011, Chapter 3)). In particular, it depends on whether or not the graph $H$ is balanced:
@@
(Janson et al., 2011, Equation (3.6)) For a fixed graph $H=(V(H),E(H))$, define
@@
where the maximum is taken over all possible non-empty subgraphs of $H.$ If the maximum is attained by $H$ itself, that is, $m(H)=\frac{|E(H)|}{|V(H)|}$, then $H$ said to be a balanced graph. Otherwise, $H$ said to be an unbalanced graph. Moreover, if the maximum in (2.1) is uniquely attained by $H$, then $H$ is called a strictly balanced graph.
@@
It is well-known from classical results (Erdős et al., 1960, Bollobás, 1981) that for an Erdős-Rényi graph random $G(n,\rho_{n})$, the threshold for the appearance of a fixed graph $H=(V(H),E(H))$ is determined by the condition $n\rho_{n}^{m(H)}\asymp 1$ (see Janson et al. (2011, Theorem 3.4)). The proof of this result can be easily adapted to show that the threshold remains the same for the inhomogeneous random graph model $G(n,\rho_{n},W)$ as well. We summarize this result in the following proposition. A proof is included in Section 3 for the sake of completeness.
@@
$X_{n}(H,G_{n})\stackrel{{\scriptstyle P}}{{\rightarrow}}\left\{\begin{array}[]{ll}0&\text{if }n\rho_{n}^{m(H)}\ll 1,\\ \infty&\text{if }n\rho_{n}^{m(H)}\gg 1.\end{array}\right$.
@@ 2.2. Statements of the Results
$Z_{n}(H,G_{n})=\frac{X_{n}(H,G_{n})-\mathbb{E}[X_{n}(H,G_{n})]}{\sqrt{\mathrm{Var}[X_{n}(H,G_{n})]}}=\frac{\Delta(H,G_{n})}{\sqrt{\mathrm{Var}[\Delta(H,G_{n})]}},$ (2.2)
where $\Delta(H,G_{n}):=X_{n}(H,G_{n})-\mathbb{E}[X_{n}(H,G_{n})]$. Recalling the the decomposition in (2.11), note that
@@
A graphon $W$ is said to be $H$-regular if, for almost every $x\in[0,1]$,
@@
for all $a\in V(H)$. Hence, if $\overline{t}(x,H,W)$ is a constant almost everywhere, then the constant must be $t(H,W)$. In other words, a graphon $W$ is $H$-regular if the homomorphism density of $H$ in $W$ when one of the vertices of $H$ is marked, is a constant independent of the value of the marked vertex. Alternatively, $H$-regularity is equivalent to the condition that the scaled $U$-statistic $\Delta_{2}(H,G_{n})$ is first-order degenerate.
@@
Hence, the notion of $K_{2}$-regularity coincides with the standard notion of degree regularity, where the degree function $d_{W}(x)$ is constant almost everywhere.
In the dense regime (where $\rho_{n}\asymp 1$), Bhattacharya et al. (2023), Hladkỳ et al. (2021) showed that $H$-regularity can result $X_{n}(H,G_{n})$ to have non-Gaussian fluctuations. The situation, however, is very different in the sparse regime. In this case, $H$-regularity implies that $\Delta_{2}(H,G_{n})$ has asymptotically negligible contribution to the total variance of $X_{n}(H,G_{n})$. Consequently, the asymptotic distribution of $X_{n}(H,G_{n})$ is determined entirely by the fluctuations of $\Delta_{1}(H,G_{n})$.
@@
Suppose Assumption 2.3 holds and the graphon $W$ is $H$-regular. Then,
@@
$Z_{n}(H,G_{n})=\frac{\Delta(H,G_{n})}{\sqrt{\mathrm{Var}[\Delta(H,G_{n})]}}\stackrel{{\scriptstyle D}}{{\rightarrow}}N(0,1),$ (2.6)
@@
Next, we consider the case where $W$ is not $H$-regular. In this case, the situation is more delicate. Specifically, an additional threshold on $\rho_{n}$ emerges, which determines whether or not $\Delta_{2}(H,G_{n})$ contributes to the fluctuations of $X_{n}(H,G_{n})$. To describe this threshold we need the following definition:
@@
(Janson et al., 2011, Equation (3.17)) For a fixed graph $H=(V(H),E(H))$ define
@@
Cycles, trees and complete graphs are strongly balanced. Figure 2 shows examples of graphs $H$ that are not strongly balanced. In both cases, for the subgraph $F$ marked by the red dotted circle, $\frac{|E(F)|}{|V(F)|-1}>\frac{|E(H)|}{|V(H)|-1}$.
@@
The notion of a strongly balanced graph was introduced by Ruciński and Vince (1986), where it is shown that $n\rho_{n}^{m_{1}(H)}\asymp 1$ is the threshold for the containment of a rooted copy of $H$ in the Erdős-Rényi model $G(n,\rho_{n})$. Recently, Maugis (2024) showed that for the sparse graphon model the number of rooted copies of $H$ (appropriately standardized) has Gaussian fluctuations, whenever $n\rho_{n}^{m_{1}(H)}\gg 1$. In Theorem 2.11 we prove that when $W$ is not $H$-regular, then $n\rho_{n}^{m_{1}(H)}\asymp 1$ marks the threshold where the asymptotic behavior of $\Delta_{1}(H,G_{n})$ and $\Delta_{2}(H,G_{n})$ undergoes a phase transition.